Solve: What will be the final volume of a 5.00 L He gas which contains 0.965 mole at 30°C and 1.00 atmosphere, if the amount of this gas is increased to 1.80 moles provided that the temperature and pressure remains unchanged?

Given: V1 = 5.00L                                                       V2 = ?

T𝟏 = 0.965 mole                                            T𝟐 = 1.80 moles



1
Expert's answer
2022-05-02T17:02:09-0400

Given:

P1 = P2 = 1.00 atm = constant

T1 = T2 = 30°C = constant

n1 = 0.965 mol

V1 = 5.00 L

n2 = 1.80 mol

V2 = unknown


Formula: V1 / n1 = V2 / n2


Solution:

Since the temperature and pressure remain unchanged, Avogadro's law can be used.

Avogadro's law can be expressed as: V1 / n1 = V2 / n2

V1n2 = V2n1

To find the final volume, solve the equation for V2:

V2 = V1n/ n1

V2 = (5.00 L × 1.80 mol) / (0.965 mol) = 9.3264 atm = 9.33 atm

V2 = 9.33 atm


Answer: The final volume of He gas will be 9.33 atm